OpenStudy (anonymous):

Use shells method to find the volum of the solid below the surface of revolution and above the xy-plane:The curve z=sin(y) (0<=y<=pi) in the yz-plane is revolved around the z-axis

OpenStudy (anonymous):

\[ \int^{\pi}_0 2\pi x\sin(x) dx = 2\pi^2 \]

OpenStudy (anonymous):

Hey man if ur struggling with this topic theres a series on youtube by a guy called derrick owens ... just search cylindrical shells and look for his name ... he Has a 9 part series and it IS EXCELLENT ... i had to teach myself this and i nailed it with his vids